Consider the line is y=3x-6
Consider the line .

Find the equation of the line that is parallel to this line and passes through the point .(2,-2)
Find the equation of the line that is perpendicular to this line and passes through the point .(2,-2)

Answer by Greenfinch(383)

The line has a gradient of 3 and a y-intercept of -6
The || line is also gradient 3. To go through (2,-2) substitute 2 and -2 for x and y so -2 = 2 + c, so c = -4
Equation is then y = 3x - 4
The perpendicular gradient is -1/3 since the two gradients multiply to -1.
So the equation is y = c - 1/3x or 3y + x = c
To determine c, substitute the values 2 and -2
-2 = 2/3 + c so c = - 2 2/3